Fast food. If four out of five doctors prefer fast food,
then at a convention of 445 doctors, how many prefer fast
food?
For this problem I subtracted one from every five and received the answer 111.

Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

Fast food. If four out of five doctors prefer fast food,
then at a convention of 445 doctors, how many prefer fast
food?
For this problem I subtracted one from every five and received the answer 111.
---------------------------
I don't see how you got 111.
4 out of 5 is 4/5 = 0.8
0.8*445 = 356

Answer by checkley77(12844)

4 out of 5 is:
4/5=.8 or 80% prefer fast food.
.8*445=356 ans. for the number of doctors who prefer fast food.